OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. 2x^4(4x^2 + 3x + 1)

OpenStudy (ash2326):

We have \[2x^4\times(4x^2+3x+1)\] Multiply \(2x^4\) to each of the terms inside the bracket and post what do you get?

OpenStudy (ash2326):

@rebecca1233 are you trying?

OpenStudy (anonymous):

yes im lost can you work me through it please

OpenStudy (ash2326):

Ok:) I'll show you how to do it \[2x^4(4x^2 + 3x + 1)\] \[2x^4 \times 4x^2+ 2x^4 \times 3x+ 2x^4 \times 1\] Multiply the constants with the constants and variables with variables \[\large 2 \times 4 x^4 \times x^2+ 2\times 3 x^4 \times x+ 2\times 1 x^4\] I'll show you first the constants' multiplication and then variables'\[\large 8 x^4 \times x^2+ 6 x^4 \times x+ 2x^4\] Now variables \[\large 8 x^{4+2}+ 6 x^{4+ 1}+ 2x^4\] Finally we get \[\large 8 x^{6}+ 6 x^{5}+ 2x^4\] Read each step and let me know if you get any doubt at any step(s):D